OpenStudy (calculusxy):

Sam takes 1/3 of an hour to walk 1 km. He continues to walk at the same rate. How many km will he walk in 4/5 hour?

OpenStudy (calculusxy):

@whpalmer4

OpenStudy (calculusxy):

@wio

OpenStudy (calculusxy):

Is it 2 2/5 km?

OpenStudy (whpalmer4):

he walks 1 km in 1/3 of an hour. \[\frac{1 km}{1/3 hr} = 3 km/hr\]right?

OpenStudy (calculusxy):

maybe

OpenStudy (calculusxy):

i am not sure.

OpenStudy (whpalmer4):

Maybe? 1 divided by 1/3 is how much?

OpenStudy (whpalmer4):

If you have a pizza and you cut it into pieces each 1/3 of the pizza, how many pieces do you get?

OpenStudy (calculusxy):

can you tell me if i am correct or not? if i am wrong can you just explain to get the correct answer?

OpenStudy (calculusxy):

i get 1/3 of the pizza right? after all , it is cut into 1/3 pieces. but should depend upon how many slices i eat :)

OpenStudy (anonymous):

do 3 divided by 5

OpenStudy (whpalmer4):

no, my point is 1 / (1/3) = 3. You get 3 slices, each slice being 1/3 of a pizza. to divide by a fraction, invert the fraction and multiply instead. 1 / (1/3) = 1* (3/1) = 3

OpenStudy (whpalmer4):

so, 3 km/hr, walking for 4/5 of an hour, 3*4/5 = 12/5 = 2 2/5 km yes, your answer is correct, but I was hoping that showing you the working would make you sufficiently confident in your own abilities that you wouldn't need to ask someone else to check it for you.

OpenStudy (calculusxy):

4/5 divided by 1/3 => 4/5 times 3/1 = 2 2/5

OpenStudy (calculusxy):

tats my wrk

OpenStudy (whpalmer4):

okay. it obscures the finding of the unit rate a bit for my tastes, but it does produce the correct answer.